OpenStudy (anonymous):

Suppose f(x)=-3x+2 and g(x)=1/2x-1. Find each value. 1.g(-2)/f(3) 2. f(-1)/g(5)

OpenStudy (anonymous):

1.g(-2)=1/2(-2)-1=-2 f(3)=-3(3)+2=-7 g(-2)/f(3)=-2/-7=2/7

OpenStudy (anonymous):

what about #2

OpenStudy (anonymous):

number two just like the steps for above. please dont be shamed to try by ur self